Differentiate with respect to π‘₯

a. (7π‘₯ – 4 )3

b. √(6π‘₯+4)


1
Expert's answer
2022-06-02T06:24:21-0400

a. "\big((7x-4)^3\big)^\prime =3(7x-4)^2\cdot 7=21(7x-4)^2"

b. "\big(\sqrt{6x+4}\big)^\prime =\big((6x+4)^{1/2}\big)^\prime=\tfrac{1}{2}(6x+4)^{-1/2}\cdot 6=\frac{3}{\sqrt{6x+4}}"


Answers: a. "21(7x-4)^2;" b. "\frac{3}{\sqrt{6x+4}}."